### v1

# Generation of near GeV protons by tightly focused laser interacting with down-ramp density plasma

Abstract
Enhancing proton energy is of great importance in laser-driven proton acceleration with finite laser energy for applications such as cancer therapy. We demonstrate an unusual acceleration scheme that achieves higher proton energies at lower laser energies by reducing the focal spot size. Through particle-in-cell simulations and theoretical modeling, we find that at small spot sizes (0.8 m), the proton energy is enhanced by 83.5%, much higher than that under conventional spot sizes (3 m). This is because the proton acceleration is dominated by electrons driven by an enhanced ponderomotive force at small spot sizes, generating stronger charge-separation fields that propagate faster. To further improve the proton energy, we analytically derive an optimal electron density profile, which enables phase-stable proton acceleration with an energy increased by 60%. These results are robust across parameter variations, suggesting that advanced focusing techniques and optimal plasma profiles could loose the requirement of laser energy, potentially reducing the dependence on large-scale laser facilities for medical and scientific applications.
I Introduction
Laser-driven plasma acceleration has emerged as a novel acceleration scheme in recent decades 1; 2; 3; 4, which has attracted extensive attention across disciplines, demonstrating broad application prospects in both scientific research and practical fields, including nuclear physics nuclear; nuclear2, quantum electrodynamics Reichwein_2025,materials science material; material2, radiography radiography; radiography2 and radiotherapy therapy; therapy0; PhysRevResearch.5.L012038. During the past 20 years, laser-driven proton acceleration has experienced rapid development, with a series of acceleration mechanisms being proposed, such as Target Normal Sheath Acceleration (TNSA) tnsa1; tnsa2, Radiation Pressure Acceleration (RPA) rpa1; rpa2; PhysRevLett.105.065002, Collisionless Shock Acceleration (CSA) PhysRevLett.101.164802; PhysRevLett.119.164801 and Magnetic Vortex Acceleration (MVA) bulanov2005ion; bulanov2010generation; bulanov2022advanced. With current laser technology, proton energies up to 150 MeV have been achieved through the interaction between intense lasers and solid targets 150mev; shou2025laser, and Au ions can reach an energy of over 1 GeV 187; martin2024narrow. For advanced acceleration mechanisms, protons have been accelerated to over 90 MeV by a novel scheme associated with sheet crossing shou2025proton; gong2020proton. The proton energy can reach the GeV level with a laser intensity of 1022 W/cm2 PhysRevLett.102.145002; PhysRevLett.103.024801; bulanov2010generation. Moreover, laser driven wakefield proton acceleration is a promising pathway for TeV proton generationyi2013proton; fp. Recently, accelerations by lasers interacting with near-critical-density (NCD) plasma targets ncd1; ncd2; ncd3 boost as a promising alternative for further energy enhancement.
When an ultra-intense laser interacts with NCD plasma, the electron mass increases due to relativistic motion, leading to a reduction in the plasma frequency. Once the plasma frequency falls below the laser frequency, the laser can penetrate the target, known as relativistic induced transparency (RIT) rit1; rit2; rit3. During this penetration, electrons experience enhanced heating via mechanisms such as direct laser acceleration (DLA) dla1; dla2, which finally improves proton energy. Additionally, in interactions with solid-density nanoscale targets, the rising edge of the laser can induce the target expansion, which results in the electron density below , leading to laser penetration because of RIT and subsequent proton energy enhancement.
Despite significant breakthroughs in acceleration mechanisms, existing laser intensities remain insufficient to accelerate protons beyond 200 MeV, a threshold required for applications such as cancer therapy therapy2; therapy3. Considering that higher laser energies generally yield higher proton energies scaling; scaling2; PhysRevE.104.025210, a critical challenge lies in maximizing proton acceleration efficiency under currently limited laser energy. Thus, optimizing laser and plasma parameters to enhance proton energies is of great importance. With the development of the wavefront correction technique, near-diffraction-limited focusing with a spot size of 1.1 m has been obtained in experiments focus. However, it is generally believed that under a certain laser intensity, a larger laser focal spot with higher laser energy results in a higher energy of the accelerated protons.
In this paper, we demonstrate that tightly focused lasers enable higher proton energies even under the same laser intensity in laser-NCD plasma interactions, which means lower laser energies. This enhancement arises because the laser ponderomotive force, which scales inversely with the square of the laser spot size, becomes significantly stronger with a small focal spot, driving more efficient electron acceleration. Consequently, a stronger and faster accelerating electric field is generated, leading to higher proton energy. Additionally, we designed a down-ramp density profile to achieve velocity matching between the protons and the accelerating electric field, further enhancing the proton energy.
II Results
We perform a series of two-dimensional particle-in-cell (PIC) simulations using the EPOCH epoch code. A Gaussian laser pulse with a normalized amplitude of , a pulse duration of 30 fs and a wavelength of 800 nm is incident normally on a fully ionized hydrogen target with a density of 20 and a thickness of 7.5 m, where is the classical critical density. The simulation domain is a 40 m × 10 m rectangular box divided into 4000 × 1000 cells, with 20 macro-particles per cell per species. To investigate the effect of laser spot size, we vary the full width at half maximum (FWHM) of the laser focal spot from 0.8 m to 5 m, with representative cases of 0.8 m and 3 m selected for detailed analysis.
II.1 Tightly focused laser
Figure 1(a) shows the temporal evolution of the maximum proton energy for the 0.8 m and 3 m spot sizes, and the proton energy spectrum at the end of acceleration is shown by the solid curve in Fig. 1(b). It can be observed that when the focal spot is 0.8 m, the cutoff energy of protons increases by 83.5% compared to the case with a 3 m spot, reaching 444 MeV. Although consuming a higher laser energy, the proton energy only reaches 242 MeV in the case with larger focal spot. This counterintuitive result indicates that reducing the laser spot size, even at a fixed intensity and lower total laser energy, enhances proton acceleration significantly. To elucidate the underlying physics, we analyze the electron acceleration dynamics for the 0.8 m case in Fig. 1(c) and the 3 m case in Fig. 1(d). When the laser interacts with the NCD plasma, the ponderomotive force expels electrons forward and laterally, forming a density spike at the laser front, which can be seen in the axial electron density distribution near x = 1 m in Figs. 1 (c) and 1 (d). In the case with a 0.8 m focal spot, a higher peak electron density can be observed, indicating stronger laser ponderomotive pushing on electrons. This results in more significant electron pile-up and more intense electron heating, as further evidenced by the electron energy spectrum represented by the dashed curve in Fig. 1(b).
To explain this observation, we propose a theoretical model. The transverse laser electric field follows a Gaussian distribution
where , is the spot size and r is the transverse distance. The longitudinal ponderomotive force fp
reaches its maximum value
at . The results clearly demonstrate that as the laser spot size decreases, the ponderomotive force, which scales inversely with the square of the spot size, becomes significantly stronger. This enhanced force drives a larger amount of electrons, consistent with the observation that in Fig. 1(d) (0.8 m spot) a higher density of ponderomotively accelerated electrons is observed compared to Fig. 1(c) (3 m spot).
To elucidate the proton acceleration mechanisms under different focal spot sizes, we present the accelerating electric field distributions and proton phase-space distributions at t = 40 and t = 60 in Figs. 2(a) and 2(b), respectively, where is the laser period. At t = 40, the laser has not yet penetrated the target and the accumulated electrons driven by the laser ponderomotive force generate a charge-separation field, referred to as the Hole-Boring hb1; hb2; hb3; hb4 field. For the 0.8 m case, a smaller focal spot results in a stronger longitudinal ponderomotive force. Here, electrons are longitudinally pushed before being transversely expelled, leading to enhanced pile-up and a larger charge-separation field, as shown in Fig. 2(a). Additionally, a stronger ponderomotive force induces a higher electron energy, causing the electron peak to propagate faster. This is evidenced by the greater displacement of the electron density peak in Fig. 1(d) (smaller spot) compared to Fig. 1(c) (larger spot) at the same time. Consequently, the electric field in Fig. 2(a) exhibits faster propagation and higher amplitude for the smaller spot, which improves the proton acceleration efficiency. At this stage, the TNSA field begins to form on the target rear side, as visible in Fig. 2(a). At t = 60, the Hole-Boring field merges with the TNSA field, forming a distinct "peak-plateau" electric field structure [Fig. 2(b)]. Protons pre-accelerated by the Hole-Boring field near the target front are then injected into the TNSA field, gaining more energy through subsequent drift acceleration.
To further clarify the proton energy disparity, Figs. 2(c) and 2(d) illustrate the proton acceleration process in the drift field, with green lines marking the trajectory of the highest-energy proton. Notably, protons are captured earlier by the accelerating field with a smaller spot [Fig. 2(d)]. Moreover, as these protons gain higher energy during the laser penetration stage, they are injected into the TNSA field before its amplitude decays, thus extracting more energy from the sheath field and achieving higher final energies. However, this abnormal increase of proton energy only occurs with small spot sizes. In the regime of a focal spot larger than 3 m, the effect of the ponderomotive force becomes insignificant because the ponderomotive force is inversely proportional to the focal spot. In this scenario, proton acceleration is primarily determined by the laser power. As a result, with increasing laser energy, electron heating becomes more intense. Consequently, under the same laser intensity with a larger focal spot of over 3 m, the proton energy tends to be higher when the focal spot is larger (Fig. 3).
To validate the robustness of these findings, we performed parameter scans of spot sizes under varying target densities and laser intensities, as shown in Fig. 3. The results universally show enhanced proton energies with reduced spot sizes. Furthermore, we propose a theoretical model where the accelerating field comprises contributions from two distinct electron groups: electrons accelerated by the ponderomotive force at the laser front and electrons accelerated by DLA mechanism in the channel. The field contributed by ponderomotive-force accelerated electrons balances the ponderomotive force
where and can be obtained from simulations. The field contributed by electrons accelerated by DLA mechanism can be solved from the Gauss’s law:
where d is the target thickness and can be obtained from simulations. Then we have:
Assuming an acceleration length of , the proton energy can be obtained as . This dual-component framework quantitatively explains the spot-size dependence of proton acceleration under various parameters. The theoretical predictions of our model are depicted by solid lines in Fig. 3, which is consistent with simulation results (scatters) in the variation trend.
II.2 Down-ramp density target
Except for laser parameter optimization, proton energies can be further increased through target design, such as employing a density down-ramp nakamura2010high; wan2019two; reichwein2021robustness, a technique widely adopted in wakefield electron acceleration to optimize injection downramp1; downramp2; downramp3, which can be achieved by the prepulse or an extra laser downramp. During the interaction of lasers with near-critical-density targets, phase-stable proton acceleration can be achieved by matching the proton velocity to the accelerating electric field. For linearly polarized lasers, the electric field velocity in Hole-Boring mechanism is given by velocity
where is the normalized laser amplitude as a replace for . The electric field amplitude satisfies:
The proton momentum needs to satisfy
to achieve velocity matching, where and . Solving this we can obtain:
The electron density profile derived from this equation is plotted as the blue curve in Fig. 4(a). An approximate solution is
shown as the red curve, where . For fixed areal density and laser parameters, protons with cutoff energies exceeding 700 MeV, a 60% increase over conventional targets, are generated with this down-ramp behind a 1.5-m-thick target, as evidenced by the spectrum in Fig. 4(b).
Figures 4(c) and 4(d) illustrate the acceleration process with the proton velocity matching the accelerating field. Protons at the field front are reflected by a shock, then exceed the field and leave the acceleration region. These protons are unable to experience the electric field and thus no longer gain energy. Meanwhile, the field’s rising velocity allows it to catch up with these protons, enabling further proton acceleration. Protons trailing the field experience stronger fields, continuously overtaking protons in front. This results in a distinct quasi-monoenergetic proton peak (50% energy spread) in Fig. 4(b), in contrast to the exponential spectrum in TNSA. As the field drifts, its velocity increases while amplitude decays [Fig. 4(d)], maintaining proton-field velocity matching throughout the acceleration progress.
III Conclusion
In summary, we demonstrate that reducing the laser spot size enables higher proton energies at lower total laser energies under a certain laser intensity. This enhancement arises from the dominant role of electrons accelerated by the ponderomotive force at small spot sizes, where the inverse-square dependence of the ponderomotive force on spot size significantly increases the electron energy. The resulting stronger charge-separation fields propagate faster, leading to more efficient proton acceleration. Meanwhile, we designed a target with a density down-ramp to achieve phase-stable proton acceleration within the electric field, thereby further enhancing the proton energy. These findings suggest that advancing laser-focusing techniques could allow proton acceleration to higher energies even under lower laser intensities, thereby relaxing the requirement of large-scale laser facilities. Combining with target optimization, the identified spot-size scaling provides a practical route to optimize laser-plasma acceleration for applications requiring high particle energies at manageable laser system sizes.